Find the velocity
abo Exrat cer edits Find the Velo cit
Answer #1

PE = KE

mgh = (1/2)mv^2

So,

V = sqrt (2gh)

= sqrt (2*9.8*3)

= 7.67 m/s

It is the value of velocity...